Tom and Gail form Owl Corporation with the following consideration:

Consideration Transferred
Basis to
Transferor
Fair Market
Value
Number of
Shares Issued
From Tom—
Cash $50,000 $50,000
Installment note 240,000 350,000 40
From Gail—
Inventory 60,000 50,000
Equipment 125,000 250,000
Patentable invention 15,000 300,000 60

The installment note has a face amount of $350,000 and was acquired last year from the sale of land held for investment purposes (adjusted basis of $240,000). Regarding these transactions, provide the following information:

If an amount is zero, enter "0".

i. Gail is considering an alternative to the plan as presented above. She is considering selling the inventory to an unrelated third party for $50,000 in the current year instead of contributing it to Owl. After the sale, she will transfer the $50,000 sales proceeds along with the equipment and patentable invention to Owl for 60 shares of Owl stock. Whether or not she pursues the alternative, she plans to sell her Owl stock in six years for an anticipated sales price of $700,000. In present value terms and assuming she later sells her Owl stock, determine the tax cost of (1) contributing the property as originally planned, or (2) pursuing the alternative she has identified.

Assume a discount rate of 6%. The present value factors at 6% are 1.000 for year 1 and 0.7050 for year 5. Further, assume Gail's marginal income tax rate is 32% and her capital gains rate is 15%.

If required, round your answers to the nearest dollar.

a. The present value of tax cost associated with sale of Owl stock for $700,000 is $.

b. The present value of the tax cost associated with the current sale of inventory for $50,000 and subsequent sale of Owl stock for $700,000 is $.
Answer #1

a. If Gail decides to contribute the inventory along with equipment and Patentable invention to Owl Corporation in exchange of 60 stock of Owl corporation,

Due to the implication of Sec 351 exemption, no capital gain would be taxable in the hands of Gail for transfer of property and the basis of the stock would be carryover basis of all the assets contributed.

Thus basis of 60 stock of Owl Corp in the hands of Gail = $60,000 + $125,000 + $15000 = $200,000

Six years from now when the stock is sold for $700,000, capital gain would be = $700,000 - $200,000 = $500,000

Tax on capital gain = $500,000 * 15% = $75,000

Present value of the tax liability = $75000 * 0.7050 = $52875

b. If Gail decides to sale the inventory now and transfer the sale proceeds to Owl corporation in exchange of 60 stock of Owl corporation,

Basis of the inventory to Gail = $60,000

Sales value of the inventory = $50,000

Therefore Ordinary loss = $60,000 - $50,000 = $10,000

Tax saving on the loss now = $10,000 * 32% = $3,200

Basis of the 60 stocks of Owl Corp in the hands of Gail = Amount contributed after sale of inventory + Carry over basis of Equipment + Carryover basis of Patentable Invention = $50,000 + $125,000 + $15000 = $190,000

Six years from now when the stock is sold for $700,000, capital gain would be = $700,000 - $190,000 = $510,000

Tax on capital gain = $510,000 * 15% = $76,500

Present value of the tax liability on capital gains = $76500 * 0.7050 = $53933

Present value of the total tax amount = $53933- $3200 = $50733

Thus the alternative of selling the inventory to third party and then contributing the proceeds to Owl Corp is a better alternative.